### v1

# Analyzing Similarity Metrics for Data Selection for Language Model Pretraining

Abstract
Similarity between training examples is used to curate pretraining datasets for language models by many methods—for diversification and to select examples similar to high-quality data. However, similarity is typically measured with off-the-shelf embedding models that are generic or trained for tasks such as retrieval. This paper introduces a framework to analyze the suitability of embedding models specifically for data curation in the language model pretraining setting. We quantify the correlation between similarity in the embedding space to similarity in pretraining loss between different training examples, and how diversifying in the embedding space affects pretraining quality. We analyze a variety of embedding models in our framework, with experiments using the Pile dataset for pretraining a 1.7B parameter decoder-only language model. We find that the embedding models we consider are all useful for pretraining data curation. Moreover, a simple approach of averaging per-token embeddings proves to be surprisingly competitive with more sophisticated embedding models—likely because the latter are not designed specifically for pretraining data curation. Indeed, we believe our analysis and evaluation framework can serve as a foundation for the design of embedding models that specifically reason about similarity in pretraining datasets.
Machine Learning, ICML
1 Introduction
The recent success of general purpose language models (Brown, 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024). Such pretraining allows the model to quickly adapt to various downstream tasks with just few-shot prompting or minimal finetuning. And so, researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from larger corpora of examples scraped from various sources.
These data curation approaches have shown promising results by improving example quality and by reducing redundancy in pretraining sets. Many of these methods use notions of similarity between examples. Similarity of an example to text from known high-quality sources (such as Wikipedia) has been used as a proxy for the quality of that example (Gunasekar et al., 2023; Penedo et al., 2024). Meanwhile, methods focused on diversification (Abbas et al., 2023; Tirumala et al., 2023) make direct use of similarity metrics to identify and remove redundant examples.
In this context, similarity between examples has been measured in terms of distances in an embedding space, and these methods have typically used generic embeddings for this purpose—e.g., intermediate representations learned by language models, or those trained for semantic retrieval. But whether these embeddings are optimal for reasoning about similarity between pretraining examples—and whether there is any benefit to using with more sophisticated and potentially expensive models—remains an open question.
Figure 1: Correlation between pretraining loss and embedding distance. Each row shows a pair of examples close in the embedding space (in particular, from the same cluster when using K-Means clustering in this space) with examples in different rows being relatively far from each other (from different clusters). We find that the close pairs of examples in each row tend to have similar pretraining losses, while there is a greater variation in losses across rows. Note that the close example pairs are ”thematically” similar, but have very different content. These results are from using the average embeddings from the final layer of a small decoder-only language model.
In this paper, we introduce a framework to analyze and quantify the suitability of a given embedding model for the task of data curation for pretraining language models. We consider the question of what a similarity metric should ideally capture for reasoning about pretraining examples. In this context, we characterize how well similarity in the embedding space correlates with similar pretraining behavior—in particular, similar pretraining losses with respect to the same model state (see Fig. 1). We also quantify whether the embedding space is able to separate examples from different data sources from among a mixture of these sources, with the assumption that these sources were curated by human expertise and combined since they were found to be complementary. And last but not least, we use the embeddings in a simple diversification recipe in curating a pretraining set and measure the gain in quality of the resulting pretrained model over a large number of downstream tasks.
We conduct these experiments using the Pile (Gao et al., 2020) as our data corpus, and in the context of pretraining a 1.7B decoder-only language model with a UL2 objective (Tay et al., 2022). We evaluate a number of different embeddings models—from representations of generic language models (e.g., trained with mask-based reconstruction or mixtures of pretraining objectives) to embeddings specifically trained for retrieval or semantic equivalence.
Our results find that most embeddings bring some utility to the pretraining data curation setting—in particular, diversity-based pretraining data curation with all of the embeddings that we consider outperforms baseline training. But perhaps surprisingly, a very simple embedding approach—derived from averaging over the independent per-token embeddings from the input layer of a small language model—performs fairly competitively to more sophisticated and computationally expensive approaches.
These results also strongly suggest that embedding models for reasoning about pretraining examples should be specialized, and that the attributes that make an embedding suitable for semantic matching and retrieval may not translate to benefits for the pretraining setting. More broadly, our work establishes a framework to evaluate current embeddings, and facilitate the design of new ones, for the important task of data curation for language model pretraining.
2 Related Work
Data Curation.
Many works have studied the problem of selecting high-quality and informative subsets from a larger corpus for efficient and effective language pretraining (Albalak et al., 2024). Indeed, works have shown that data curation for pretraining (both for language and vision-language models) improves neural scaling laws (Sorscher et al., 2022), and that curation techniques should be aware of the particular compute budget (Goyal et al., 2024).
One family of works approaches data curation with the goal of retaining only high quality examples. The simplest approaches in this family look at heuristics that can filter out noisy or extremely low-quality instances such as documents containing less than 5 sentences or those with extremely short lines of text (Raffel et al., 2020; Xue, 2020). Beyond these heuristic approaches, a line of work focus on extracting high-quality examples based on similarity to known high-quality sources, such as Wikipedia or textbooks (Gunasekar et al., 2023; Penedo et al., 2024). Other works look at creating targeted sets—defining quality in terms of their relevance for particular downstream tasks (Xie et al., 2023), with some creating these by adaptively updating weights over domain mixtures (Xie et al., 2024; Jiang et al., 2024).
The other family of data curation approaches focus on pruning datasets to ensure a notion of diversity (Abbas et al., 2023, 2024; Tirumala et al., 2023). While some of these works (Abbas et al., 2024; Tirumala et al., 2023) compare different embedding models as part of brief ablations, these comparisons are in the narrow context of their specific divsersification algorithms, and do not provide intuition as to why one embedding may perform better than another.
Finally, it is worth mentioning that some works in data selection use the model being trained as part of the selection process, often through the use of influence functions or gradient-based methods (Garima et al., 2020; Xia et al., 2024; Engstrom et al., 2024). However, these are typically used during for data-curation for multi-epoch training on small datasets or only during the finetuning stage, and would be prohibitively expensive to use for pretraining.
Text Embedding Models.
Another related line of work is the task of learning embedding models for text. Many various approaches to learn text embedding models, with objectives including mask-based reconstruction (Devlin, 2018; Liu et al., 2019), a combination of multiple different tasks (Cer, 2018), and contrastive learning-based approaches (Gao et al., 2021; Neelakantan et al., 2022; Izacard et al., 2022; Lee et al., 2024). More recent work has demonstrated extracting embeddings via prompting, specifically designed for predicting model performance (Sam et al., 2025).
These embedding models have largely been studied in the context for classification or similarity measures (Gomaa et al., 2013; Agirre et al., 2013, 2016), and recent focus has been improving performance on aggregate on large-scale benchmarks (Muennighoff et al., 2022) that are comprised of multiple tasks (e.g., retrieval, clustering, classification, ranking). Many models that achieve strong performance on these large-scale benchmarks have benefitted from scaling (Jiang et al., 2023; Chowdhery et al., 2023), with the help of distilled information from autoregressive models (Lee et al., 2024). However, this has conflicting incentives with their utility in the pretraining setting, as large text embedding models are impractical and too computationally expensive to run inference over the full pretraining corpus. To the best of our knowledge, our work provides the first study of various text embedding models for pretraining data curation.
3 Evaluating Embeddings for Pretraining Data Curation
We now describe our approach to analyze and evaluate a given embedding model in terms of its suitability for reasoning about similarity among pretraining examples. As mentioned above, similarity is used to (a) find examples that are similar to known “desirable” examples (e.g., examples of known high-quality or those representing downstream tasks of interest) and to (b) discover and remove redundancies in the pretraining corpus. Accordingly, we design experiments that measure an text embedding model’s performance towards these criteria.
3.1 Evaluating Correlation with Pretraining Loss
We begin by evaluating whether low distances in a model’s embedding space correlates with similar values of difficulty, which we measure by intermediate losses during language model pretraining. To do so, we first use a balanced K-Means clustering algorithm in the given embedding space to cluster all examples in the pretraining set for a target cluster size. Then, we look at the variance of loss values within each cluster. We note that we use balanced clustering (allowing only some variation in the sizes of different clusters) to ensure that average within-cluster variance is comparable across clusterings in different embedding spaces. We repeat this process for a multiple target cluster sizes for all embeddings.
When reporting our results, we measure variance reduction, or the ratio of the overall variance across all examples in the pretraining dataset to the variance in loss computed within clusters in the given embedding space. Specifically, we define variance reduction as
$V(C)$ $=\frac{E_{x\sim D}[(\ell(x)-E[\ell(x)])^{2}]}{E_{C_{i}\sim\mathbbm{C}}\left[E_{x\sim C_{i}}[(\ell(x)-E_{x\sim C_{i}}[\ell(X)])^{2}]\right]},$
where $C=\{C_{1},...,C_{m}\}$ is a partition of all datapoints in the dataset $D$, and $\mathbbm{C}$ is a uniform distribution over each disjoint set in the partition $C$. This metric can be interpreted as a measure of predicting pretraining generalization, as nearby points have similar loss values. We remark that random clusters achieve a variance reduction of 1, and larger values of variance reduction imply that points with more similar pretraining loss are clustered together.
A high value of variance reduction implies that similarity in the embedding space correlates well with pretraining loss. While an isolated pair of examples having similar pretraining loss may be entirely unrelated in quality and content, the fact that an embedding space that consistently brings together examples of similar loss values strongly suggests that these similar examples will behave similarly in terms of contributing to the quality of the language model.
Beyond finding examples of similar quality and utility, a high variance reduction score implies that the clustering above can serve as a particularly useful proxy for more dynamic and online data sampling strategies (Xie et al., 2024), reducing the number of required forward passes in strategies where datapoints are selected based on their current pretraining loss. This also has implications towards approaches that look to self-improve models through propagating labeled information to nearby examples (Wei et al., 2021; Cai et al., 2021; Pukdee et al., 2023), such as weak-to-strong generalization (Burns et al., 2023).
3.2 Measuring Cluster Purity of Data Sources
Next, we study the purity of the respective (K-Means) clusters generated with each considered embedding model with respect to the underlying data source or mixture. Large pretraining datasets such as the Pile (Gao et al., 2020) are comprised of various distinct yet complementary hand-curated data sources (e.g., high quality sources such as Wikipedia, code-based data, data from medical domains). While there may be some similarities between datapoints from different domains, we believe that clusters should generally group points from the same domain together and separate isntances from different domains. In fact, we remark that since embedding models are generally not trained with explicit knowledge of source labels or any domain metadata, and the ability to group together data from the same source and distinguish those from different sources shows an embedding model’s alignment with human judgement (or at least that of the dataset curators).
Letting $s\in\{0,\cdots,S\}$ represent the index of the source in some finite number of sources $S$, we compute the cluster purity of a set of clusters $C$ as
$P(C)$ $=\mathbbm{E}_{C_{i}\sim\mathbbm{C}}\left[\frac{\max_{s}|C_{i}\cap D_{s}|}{|C_{i}|}\right],$
where $D_{s}$ represents the subset of datapoints that are from source $s$. Intuitively, this represents the proportion of the maximum frequency source to the total cluster size. A cluster purity of 1 represents a clustering that completely separates examples from different domains.
3.3 Diversification-based Pretraining Data Curation
Prior work has demonstrated that diversification, based on similarity in embedding spaces, can yield more effective pretraining datasets that in-turn lead to improved models (Tirumala et al., 2023). While the diversification techniques themselves tend to be sophisticated, our goal here is to evaluate the utility of a given embedding for diversification. Therefore, we use a simplified version of these diversification of approaches to (1) select a subset from a larger pretraining corpus, (2) pretrain a model on this curated subset, and then (3) report performance on a large set of downstream tasks.
Figure 2: Reduction in variance in loss among clusters (e.g., overall variance / clustering variance) as increase the number of gradient steps in pretraining (left). Reduction in variance as we vary the approximate cluster size used in K-means for clustering examples with various embedding models (right). We compute our results on the checkpoint after 26k gradient steps. In both figures, we compute our results over 50 million sampled clusters from the Pile, and we remark that larger values are better. The overall pretraining loss variances for checkpoints of 26k, 50k, and 76k training steps are 0.521, 0.508, and 0.489 respectively. We find that embeddings extracted from a small version of the downstream language model being trained best cluster points with similar pretraining losses, across different intermediate training checkpoints and cluster sizes.
We begin by clustering the larger pretraining corpus in the given embedding space; in contrast to Section 3.1 and Section 3.2, we do not use balanced clustering which can lead to clusters with a wide variation in “diameters” (distances among points in the same cluster). Instead, since the goal here is diversification, we find clusters such that distances between all pairs of points in the same cluster are within a specified threshold $\epsilon$.
Given clusters of pretraining data, prior work often uses complex pruning or sampling schemes from large clusters. For simplicity, we simply take the point in each cluster that is closest to the cluster centroid (i.e., the average embedding of all points in the cluster), which likely represents the most representative example for that cluster. We note that prior work (Tirumala et al., 2023) emphasizes the importance of a deduplication step in language model pretraining, and our curation strategy implicitly performs this process when selecting only a single point from each cluster.
Given that we select a single point from each cluster, this implies that we need the number of clusters to be equal or greater to the number of desired pretraining examples. In practice, we sweep a large grid of $\epsilon$ values and select the largest threshold that still produces a sufficient number of clusters. Another important point of note is that producing such a large number of clusters is computationally expensive. To be able to scale to such a large number of clusters, we use the reciprocal agglomerative clustering (RAC) (Sumengen et al., 2021) algorithm. This has the advantage of building a clustering from progressively larger values of $\epsilon$, which works well with our need of a grid search over values of $\epsilon$.
Finally, we pretrain a language model on the selected subset of examples, and measure few-shot performance of the pretrained model on a diverse variety of downstream tasks.
4 Experiments
We analyze various different approaches to extract text embeddings across multiple different pretraining settings.
4.1 Experiment Details
Text Embedding Models
For text embedding models, we compare Universal Sentence Encoders (USE) (Cer, 2018), which are general purpose text encoders, and Gecko (Lee et al., 2024), a retrieval-focused text embedding model trained via synthetic data distilled from LLMs, and BERT (Devlin, 2018) models. We also compare against various techniques to extract embeddings from a small version of the downstream language model we are training (e.g., a  200M parameter language model). First, we extract an embedding by simply averaging the token embeddings matrix over all tokens in the input sequence, which we refer to as LM Token Embeds. We note that this is equivalent to a learned unigram model (i.e., ignoring all positional information). We also extract an embeddings from a forward pass of the language model by averaging the activations over all tokens in the input sequence at the final hidden layer, which we refer to as LM Output Embeds. In our diversity-based pretraining data curation experiments, we also add a comparison to a naive, random subset selection (Baseline) to evaluate how much benefit is added from our standardized curation strategy with these various embedding models.
To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500000 examples) and extract the top 64 dimensions upon which to project. The projected embeddings are then normalized to have a unit L2 norm. We also explore using random projections (Bingham & Mannila, 2001) as a different dimensionality reduction strategy, finding that random projections do not perform as well as PCA (see Section 4.5 for ablation studies).
Dataset Details
For all of our experiments (e.g., pretraining data curation and predicting loss generalization), we use the Pile (Gao et al., 2020). We pack together documents into a sequence of length 1280, with “[eod]” as delimeters between documents.
Pretraining Details
For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022), on the curated subset of the Pile data through the process outlined in Section 3.3. We use a selection budget of 20B tokens, which is approximately 20% of sequences from the Pile. This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
We pretrain with a learning rate of 0.001 with a linear decay and a batch size of 1024. For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens. For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown (2020). This invovles a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks. The full details about our evaluation sets are deferred to Section B.1. For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 20B randomly selected tokens.
4.2 Pretraining Loss Correlation Results
To measure the ability of embedding models to reflect pretraining loss generalization and to cluster together datapoints with similar difficulty, we report their variance reduction as we vary (1) the number of gradient updates in language model pretraining and (2) as we vary the average cluster size in our K-Means clustering.
We observe that embeddings produced via a small UL2 model (i.e., LM Token Embeds and LM Output Embeds) trained on the same data achieve a much larger reduction in loss variance when compared to off-the-shelf models; these embeddings produce better clusters with datapoints with similar intermediate pretraining losses (Figure 2). These trends hold across multiple different intermediate training checkpoints and for all different cluster sizes. This suggests that training models specifically on the data to perform data curation is preferable to using off-the-shelf embedding models for predicting pretraining loss generalization. Furthermore, we also generally observe that newer embedding models (e.g., Gecko) do not improve upon older alternatives (e.g., BERT and USE), supporting that embedding model improvements in other settings such as retrieval do not translate to this application in the pretraining setting.
Visualizing Within-Cluster Examples
To better interpret what is computed by these similarity metrics, we also visualize pairs of datapoints from the various clusters in Figure 1 to understand what types of examples are placed closely together and which have similar pretraining losses. We randomly sample 3 clusters produced by K-Means clusters with the smallest average cluster size with the LM Output Embeds method to extract embeddings. We randomly sample 2 points from these 3 clusters and observe that samples from the same cluster are thematically similar, although they contains very different information (e.g., the last example referring to two different people in history with some governmental relations, although from different times and different countries). We observe that pretraining losses are similar between within-cluster examples, and lower distances correspond to lower differences in pretraining loss.
4.3 Cluster Purity Results
Figure 3: Comparison of the purity with respect to data source of K-Means clustering produced by various embedding models on the Pile. Cluster purities are averaged over 50 million clusters.
To better understand and interpret the similarity metrics defined by these embedding models, we measure the purity of clusters with respect to the underlying data source. We again remark that none of these embedding models have been trained with knowledge of the data source. Overall, we observe that most embedding models produce fairly pure clusters, where a majority of points come from the same underlying data source (Figure 3). This supports that these embeddings models are generally aligned with human judgement about differences between types of data.
We observe that the embeddings extracted from the small version of the trained downstream language model achieve the highest cluster purity. We also note that Gecko embeddings achieve the third-highest cluster purity, whereas it performs poorly on the task of predicting loss generalization through variance reduction. This suggests that an improved ability to predict pretraining loss generalization cannot simply be explained via producing clusters that are more pure with respect to underlying data sources. Future work could explore incorporating domain information or metadata into the embedding model’s learning objective, as a form of weak supervision (Sam & Kolter, 2023).
Task Baseline USE Gecko BERT LM Token Embeds LM Output Embeds
ARC - Challenge 32.44 32.99 33.73 32.71 32.54 33.53
ARC - Easy 63.78 65.12 64.24 65.08 64.91 65.46
BoolQ 56.47 59.18 58.31 60.08 62.86 61.93
SuperGLUE - CB 42.41 43.45 48.21 41.67 42.86 48.81
SuperGLUE - Copa 75.25 76.33 77.33 74.00 74.67 78.00
HellaSwag 54.98 56.71 57.11 56.54 56.76 57.53
Multi RC 57.71 55.69 55.91 56.27 55.95 55.19
OpenBook QA 46.30 46.20 46.27 46.60 45.93 46.67
PiQA 72.27 72.98 72.74 73.58 72.76 73.63
Race H 38.02 37.94 38.47 38.75 38.84 38.70
Race M 51.71 52.00 53.32 52.88 52.09 52.60
ReCoRD 84.95 84.35 84.64 84.71 84.51 84.91
RTE 54.51 54.87 51.50 54.87 51.62 52.83
Story Cloze 73.93 74.10 74.26 73.95 73.26 74.49
WiC 48.20 47.86 47.49 47.54 48.17 47.28
Winograd 73.99 75.95 77.05 77.29 77.05 76.56
WinoGrande 59.23 58.91 59.01 59.46 59.96 59.25
WSC 74.12 73.92 73.68 74.74 73.22 74.97
Lambada 21.37 29.86 40.12 34.69 31.43 33.82
Natural Questions 10.01 10.76 10.13 10.54 11.13 11.42
SQuAD v2 51.05 54.26 51.70 51.79 58.56 55.37
TriviaQA Wiki 34.39 35.01 32.99 34.95 34.29 36.08
Web Questions 17.05 17.04 17.87 17.98 18.73 19.13
Average 51.92 52.85 53.30 53.07 53.13 53.83
Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 20B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective. Bolding and underlining denote the best and second-best performing methods on each task, respectively. Results are averaged over 3 pretraining runs. We find that embeddings extracted from small version of the downstream language model and Gecko perform the best for pretraining data curation.
4.4 Diversification-based Pretraining Data Curation Results
We report the average downstream task performance of language models trained on pretraining data subsets produced by using various embedding models in the simple diversity-based data curation strategy in Table 1. We first observe that the standardized diversity-based data curation technique applied to all embedding models (both off-the-shelf and those based on small downstream trained language models) outperform the naive random subset selection baseline. Secondly, we observe the LM Output Embeds and Gecko embeddings achieve the best and second-best performance when averaged across all tasks, respectively.
A notable result is that LM Token Embeds performs comparably to Gecko, while outperforming all the other off-the-shelf embedding models. This is of practical interest due to the simplicity of producing these embeddings; they are simply constructed via averaging over the learned token embeddings and, consequently, are extremely quick to compute in comparison to all other embedding methods. Thus, LM Token Embeds serves as a powerful alternative in settings that are compute-limited.
These pretraining results also suggest that even simple notions of similarity, even those that do not account for positional information, are sufficient for many pretraining data curation applications. Finally, they support that the training of models specifically for this task (e.g., on the same pretraining dataset) is preferable to the use of general purpose off-the-shelf embedding models.
Figure 4: Ablation on the number of components in PCA for Gecko and LM Output Embeds. Results are averaged over 50 million sampled clusters from the Pile. Using more components in PCA better cluster points with similar pretraining loss.
Figure 5: Ablation comparing the use of PCA or Random projections for dimensionality reduction in Gecko and LM Output Embeds. Results are averaged over 50 million sampled clusters from the Pile. Dimensionality reduction via PCA performs better for variance reduction when compared to Random Projections.
4.5 Ablations
We present ablations on the various dimensionality reduction strategies (i.e., technique and resulting size) that we use for the embedding models before running clustering. We present the ablations on Gecko and LM Output Embeds here and defer results on others embeddings to Appendix A.
Impact of Dimensionality on Variance Reduction
The dimensionality of the embeddings used in clustering often must be low for pretraining scales. Here, we run an ablation studying the role of dimensionality (i.e., the number of components in PCA onto which the embeddings are projected) in the variance reduction through clustering (Figure 4). We remark that scaling clustering to accommodate smaller average cluster sizes (i.e., more cluster centers) is intractable for embeddings that have high dimensionality. Thus, we can only report results on various embeddings with a large number (i.e., 256) of components in PCA with larger average cluster sizes. We observe the trend across all embeddings that a higher dimension and larger number of PCA components improves the embedding models’ ability to cluster points by pretraining loss (see other embedding model results in Appendix A.1).
Dimensionality Reduction Techniques
Another common technique to perform dimensionality reduction (especially with respect to maintaining pairwise distances and similarities) is to use random projections. We experiment with using random projections in our loss clustering experiments and observe that it is outperformed by PCA in terms of variance reduction across all cluster sizes (Figure 5).
5 Discussion
We analyze the fundamental role of computing similarity in pretraining data curation, finding that coarse-grained notions of similarity are particularly useful and that specific models (even simple ones based on averaging learned token embeddings) can outperform other off-the-shelf embedding models. We believe that both our results and the new evaluation framework for similarity metrics provide new insights into the role of similarity in pretraining data curation. For instance, our variance reduction results shows that embedding models tightly cluster points with similar values of pretraining losses, which has larger implications for dynamic data selection. Future work could study that, during training, we can compute the loss on the centroid in each cluster, and use this to select clusters on which to train, that contain points with high values of these intermediate loss. Our method has other, broad potential applications in the pretraining data setting, such as selecting task-specific data for finetuning (Xia et al., 2024) or even in evaluating the generation of synthetic data for (Meng et al., 2022; Sam et al., 2024) by looking at the distance between these potential data to be selected or generated and real data. Finally, our results highlight the benefits for designing models specifically for language model pretraining settings, and our framework provides an evaluation to exactly capture their utility.
Appendix A Additional Experimental Results
A.1 Impact of Dimensionality Reduction on Variance Reduction
We now present the remaining results for other embedding models, when we ablate the number of components used in PCA for K-Means clustering, specifically when looking at the reduction in variance of pretraining loss of points within the same cluster. We observe that more components in PCA indeed help achieve higher variance reduction across all embedding models.
Figure 6: Ablation on the number of components in PCA for embeddings from BERT, USE, and LM Token Embeds. Results are averaged over 50 million sampled clusters from the Pile.
A.2 Comparison of Random Projections to PCA for Dimensionality Reduction
We now present the remaining results for other embedding models, when we use Random Projections for dimensionality reduction instead of using PCA. We consistently observe that embeddings paired with PCA outperform those using Random Projections.
Figure 7: Ablation where we compare using Random Projections for dimensionality reduction with embeddings from BERT, USE, and LM Token Embeds. Results are averaged over 50 million sampled clusters from the Pile.
Appendix B Additional Experimental Details
B.1 Additional Evaluation Details
We report results averaged over a large numer of downstream tasks. These datasets largely follow two categories: scoring and decoding tasks. Scoring tasks primarly look at the output distribution of the model, while decoding tasks look at text generations from the language model. Scoring tasks are performed as 1-shot (i.e., giving one demonstration of format and answer), while decoding is performed zeroshot. For scoring tasks, we look at the standard top-1 accuracy. The list of scoring tasks is as follows:
  • ARC Challenge: Easy and Challenge (Clark et al., 2018)
  • BoolQ (Clark et al., 2019)
  • SuperGLUE - CB and Copa (Wang et al., 2019)
  • HellaSwag (Zellers et al., 2019)
  • MultiRC (Khashabi et al., 2018)
  • OpenBookQA (Mihaylov et al., 2018)
  • PIQA (Bisk et al., 2020)
  • RACE-H and RACE-M (Lai et al., 2017)
  • ReCoRD (Zhang et al., 2018)
  • RTE (Dagan et al., 2010)
  • Story Cloze (Mostafazadeh et al., 2016)
  • WIC (Pilehvar & Camacho-Collados, 2018)
  • Winograd (Levesque et al., 2012)
  • WinoGrande (Sakaguchi et al., 2021)
  • WSC (Levesque et al., 2012)
For decoding or text generation tasks, we evaluate the language model outputs with its F1 score. Decoder tasks are also evaluated as 1-shot. The list of decoding tasks is as follows:
  • Lambada (Paperno et al., 2016)
  • Natural Questions (Kwiatkowski et al., 2019)
  • Squad v2 (Rajpurkar et al., 2018)
  • Trivia QA Wiki (Joshi et al., 2017)
  • Web Questions (Berant et al., 2013)
B.2 Additional Embedding Model Details
The embedding models of BERT, Gecko and USE are trained with sequence lengths of 512, which we apply on the first 512 tokens of data from the Pile. For Gecko, we use the 110 million parameter model version, while for USE we use the 109 million parameter version. For BERT, we also use a 109 million parameter model. For the small language model that we train with the UL2 objective, we use one with approximately 200 million parameters.
Both the off-the-shelf BERT and USE embedding models have a dimensionality of 512. The Gecko embedding model has a dimensionality of 768. The small language model has a token embedding dimension of 512 and an hidden activation dimension of 512, which means that both LM Token Embeds and LM Output Embeds have 512 dimensions.
B.3 Additional Hyperparameter Details
Clustering
For our RAC clustering for pretraining experiments at a 20B token budget, we use a value of $\epsilon$ as the particular diameter of clusters:
  • USE: $\epsilon=0.2$, which defines roughly 225 million clusters
  • Gecko: $\epsilon=0.2$, which defines roughly 220 million clusters
  • BERT: $\epsilon=0.001$, which defines roughly 175 million clusters
  • LM Token Embeds: $\epsilon=0.001$, which defines roughly 170 million clusters
  • LM Output Embeds: $\epsilon=0.03$, which defines roughly 180 million clusters
For our K-Means clustering, we perform clustering at 4 different levels of granularity in our variance reduction and cluster purity results. We create four sets of clusterings with an average cluster size of 25, 50, 100, 150, with a minimum cluster size of $\frac{1}{5}$ times the average cluster size and a maximum cluster size of $5$ times the average cluster size. For both RAC and K-Means, we use the squared L2 (Euclidean) distance.
B.4 Dimensionality Reduction Details
For running our dimensionality reduction via PCA, we compute the means and components on which to project on a subset of the data ($\sim$500000 points). We first standardize the embeddings to have a mean of 0 and variance of 1 before running PCA. As previously mentioned, after projecting onto the desired number of principal components, we perform L2 normalization.
For our random projections, we use a sparse random projections onto values of $-\frac{\sqrt{n}}{\sqrt{64}},0,\frac{\sqrt{n}}{\sqrt{64}}$ with probabilities $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ respectively (i.e., the default parameters via scikit-learn). We also L2 normalize the result of random projections.
B.5 Compute Details
Pretraining experiments for our 1.7B parameter language models are run on 512 v5 TPUs, where each pretraining run takes approximately 3 days.

### v2

@@
 1 Introduction
-The recent success of general purpose language models (Brown, 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024).
+The recent success of general purpose language models (Brown et al., 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024).
 Such pretraining allows the model to quickly adapt to various downstream tasks with just few-shot prompting or minimal finetuning.
@@
 In our diversity-based pretraining data curation experiments, we also add a comparison to a naive, random subset selection (Baseline) to evaluate how much benefit is added from our standardized curation strategy with these various embedding models.
-To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500000 examples) and extract the top 64 dimensions upon which to project.
+To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500,000 examples) and extract the top 64 dimensions upon which to project.
 The projected embeddings are then normalized to have a unit L2 norm.
@@
 For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022), on the curated subset of the Pile data through the process outlined in Section 3.3.
-We use a selection budget of 20B tokens, which is approximately 20% of sequences from the Pile.
+We use a selection budget of 200B tokens, which is approximately 20% of sequences from the Pile.
 This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
@@
 For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens.
-For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown (2020).
+For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown et al.
+(2020).
 This invovles a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks.
 The full details about our evaluation sets are deferred to Section B.1.
-For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 20B randomly selected tokens.
+For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 200B randomly selected tokens.
 4.2 Pretraining Loss Correlation Results
@@
 Average 51.92 52.85 53.30 53.07 53.13 53.83
-Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 20B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective.
+Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 200B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective.
 Bolding and underlining denote the best and second-best performing methods on each task, respectively.
@@
 Clustering
-For our RAC clustering for pretraining experiments at a 20B token budget, we use a value of $\epsilon$ as the particular diameter of clusters:
+For performing RAC clustering for our pretraining experiments, we use a value of $\epsilon$ as the particular diameter of clusters:
   • USE: $\epsilon=0.2$, which defines roughly 225 million clusters
@@
 B.4 Dimensionality Reduction Details
-For running our dimensionality reduction via PCA, we compute the means and components on which to project on a subset of the data ($\sim$500000 points).
+For running our dimensionality reduction via PCA, we compute the means and components on which to project on a subset of the data ($\sim$500,000 points).
 We first standardize the embeddings to have a mean of 0 and variance of 1 before running PCA.

### v3

@@
 Abstract
-Similarity between training examples is used to curate pretraining datasets for language models by many methods—for diversification and to select examples similar to high-quality data.
-However, similarity is typically measured with off-the-shelf embedding models that are generic or trained for tasks such as retrieval.
-This paper introduces a framework to analyze the suitability of embedding models specifically for data curation in the language model pretraining setting.
-We quantify the correlation between similarity in the embedding space to similarity in pretraining loss between different training examples, and how diversifying in the embedding space affects pretraining quality.
-We analyze a variety of embedding models in our framework, with experiments using the Pile dataset for pretraining a 1.7B parameter decoder-only language model.
-We find that the embedding models we consider are all useful for pretraining data curation.
-Moreover, a simple approach of averaging per-token embeddings proves to be surprisingly competitive with more sophisticated embedding models—likely because the latter are not designed specifically for pretraining data curation.
-Indeed, we believe our analysis and evaluation framework can serve as a foundation for the design of embedding models that specifically reason about similarity in pretraining datasets.
-Machine Learning, ICML
+Measuring similarity between training examples is critical for curating high-quality and diverse pretraining datasets for language models.
+However, similarity is typically computed with a generic off-the-shelf embedding model that has been trained for tasks such as retrieval.
+Whether these embedding-based similarity metrics are well-suited for pretraining data selection remains largely unexplored.
+In this paper, we propose a new framework to assess the suitability of a similarity metric specifically for data curation in language model pretraining applications.
+Our framework’s first evaluation criterion captures how well distances reflect generalization in pretraining loss between different training examples.
+Next, we use each embedding model to guide a standard diversity-based data curation algorithm and measure its utility by pretraining a language model on the selected data and evaluating downstream task performance.
+Finally, we evaluate the capabilities of embeddings to distinguish between examples from different data sources.
+With these evaluations, we demonstrate that standard off-the-shelf embedding models are not well-suited for the pretraining data curation setting, underperforming even remarkably simple embeddings that are extracted from models trained on the same pretraining corpus.
+Our experiments are performed on the Pile, for pretraining a 1.7B parameter language model on 200B tokens.
+We believe our analysis and evaluation framework serves as a foundation for the future design of embeddings that specifically reason about similarity in pretraining datasets.
 1 Introduction
-The recent success of general purpose language models (Brown et al., 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024).
-Such pretraining allows the model to quickly adapt to various downstream tasks with just few-shot prompting or minimal finetuning.
-And so, researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from larger corpora of examples scraped from various sources.
-These data curation approaches have shown promising results by improving example quality and by reducing redundancy in pretraining sets.
+The recent success of language models (Brown et al., 2020; Chowdhery et al., 2023) is in no small part due to pretraining on large and diverse text corpora scraped from a variety of sources (Raffel et al., 2020; Gao et al., 2020; Penedo et al., 2024).
+Researchers have explored a variety of approaches to assemble effective pretraining sets, typically by selecting a high-quality and diverse subset from a larger corpus of examples scraped from multiple data sources.
+These data curation approaches have shown promising results by improving example quality and reducing redundancy in pretraining sets.
 Many of these methods use notions of similarity between examples.
@@
 Meanwhile, methods focused on diversification (Abbas et al., 2023; Tirumala et al., 2023) make direct use of similarity metrics to identify and remove redundant examples.
-In this context, similarity between examples has been measured in terms of distances in an embedding space, and these methods have typically used generic embeddings for this purpose—e.g., intermediate representations learned by language models, or those trained for semantic retrieval.
-But whether these embeddings are optimal for reasoning about similarity between pretraining examples—and whether there is any benefit to using with more sophisticated and potentially expensive models—remains an open question.
-Figure 1: Correlation between pretraining loss and embedding distance.
-Each row shows a pair of examples close in the embedding space (in particular, from the same cluster when using K-Means clustering in this space) with examples in different rows being relatively far from each other (from different clusters).
-We find that the close pairs of examples in each row tend to have similar pretraining losses, while there is a greater variation in losses across rows.
-Note that the close example pairs are ”thematically” similar, but have very different content.
-These results are from using the average embeddings from the final layer of a small decoder-only language model.
-In this paper, we introduce a framework to analyze and quantify the suitability of a given embedding model for the task of data curation for pretraining language models.
-We consider the question of what a similarity metric should ideally capture for reasoning about pretraining examples.
-In this context, we characterize how well similarity in the embedding space correlates with similar pretraining behavior—in particular, similar pretraining losses with respect to the same model state (see Fig.
+In this context, similarity between training examples has often been measured in terms of distances in an embedding space.
+Many approaches have typically used generic embeddings for this purpose (Chang and Jia, 2023; Vo et al., )—off-the-shelf embedding models that have been trained for tasks such as semantic retrieval or mask-based reconstruction.
+However, whether these embeddings are optimal for reasoning about similarity between pretraining examples and whether there is any benefit to using more sophisticated and computationally expensive models remains an open question.
+Furthermore, evaluating such embeddings in pretraining applications is fundamentally challenging, as the sheer scale of pretraining data makes many experiments intractable.
+Such challenges necessitate specialized evaluations — new criteria for assessing the utility of embeddings in pretraining applications.
+In this paper, we propose a novel evaluation framework to assess the suitability of an embedding model for curating pretraining data for language models.
+Our goal is to establish a new standard for evaluating similarity metrics in the context of language model pretraining.
+We begin by asking: what should a similarity metric ideally capture to be useful in this setting?
+Unlike prior work that evaluates embeddings for downstream tasks like retrieval or classification, we argue that embeddings used for data curation must reflect training dynamics—such as generalization behavior or corpus redundancy—that are unique to the pretraining setting.
+Figure 1: A visualization of the correlation between pretraining loss and embedding distance.
+Each row shows a pair of examples close in embedding space (from the same K-means cluster), with examples in different rows being far from each other (from different clusters).
+We find that close pairs of examples tend to have similar pretraining losses, while there is a greater variation in losses across clusters.
+Close example pairs are "thematically" similar but have different content.
+These results are from averaged embeddings from the final layer of a small decoder-only language model.
+Our first evaluation criterion measures how well distances in the embedding space correlate with generalization behavior during pretraining—specifically, whether nearby examples in embedding space exhibit similar pretraining loss under the same model state (see Fig.
 1).
-We also quantify whether the embedding space is able to separate examples from different data sources from among a mixture of these sources, with the assumption that these sources were curated by human expertise and combined since they were found to be complementary.
-And last but not least, we use the embeddings in a simple diversification recipe in curating a pretraining set and measure the gain in quality of the resulting pretrained model over a large number of downstream tasks.
-We conduct these experiments using the Pile (Gao et al., 2020) as our data corpus, and in the context of pretraining a 1.7B decoder-only language model with a UL2 objective (Tay et al., 2022).
-We evaluate a number of different embeddings models—from representations of generic language models (e.g., trained with mask-based reconstruction or mixtures of pretraining objectives) to embeddings specifically trained for retrieval or semantic equivalence.
-Our results find that most embeddings bring some utility to the pretraining data curation setting—in particular, diversity-based pretraining data curation with all of the embeddings that we consider outperforms baseline training.
-But perhaps surprisingly, a very simple embedding approach—derived from averaging over the independent per-token embeddings from the input layer of a small language model—performs fairly competitively to more sophisticated and computationally expensive approaches.
-These results also strongly suggest that embedding models for reasoning about pretraining examples should be specialized, and that the attributes that make an embedding suitable for semantic matching and retrieval may not translate to benefits for the pretraining setting.
-More broadly, our work establishes a framework to evaluate current embeddings, and facilitate the design of new ones, for the important task of data curation for language model pretraining.
+This captures their utility of similarity metrics in loss-based data selection strategies (Jiang et al., 2023) or for selecting examples similar to high-quality references (Penedo et al., 2024).
+Next, we evaluate the embedding model’s utility in guiding a simple diversity-based curation method (similar to those employed in (Tirumala et al., 2023)) and assess its effectiveness by pretraining a model on the curated subset and measuring downstream task performance.
+Finally, we also measure whether the embedding space can distinguish between examples from different data sources.
+While not directly used in data selection, this serves as a useful proxy for embedding quality under the assumption that these sources were curated by human expertise and reflect meaningful human-curated structure.
+We conduct all of our experiments using the Pile (Gao et al., 2020) as our data corpus, and in the context of pretraining a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022) on 200B tokens.
+We evaluate a number of different off-the-shelf embedding models—from representations of generic language models (e.g., trained with mask-based reconstruction or mixtures of pretraining objectives) to embeddings specifically trained for retrieval or semantic equivalence.
+We compare these to two specialized approaches that are derived from smaller, proxy models of the downstream model being trained on the same pretraining set: (1) computing the average token embedding at the final hidden state, and (2) a significantly more computationally efficient and simple average of the token-embeddings from the input layer (i.e., requiring no forward passes).
+Despite their strong performance on standard semantic similarity and retrieval benchmarks, we find that off-the-shelf embedding models underperform in the context of pretraining data curation, suggesting that existing benchmarks may not reflect the inductive biases most relevant to language model pretraining.
+Surprisingly, even this simple method that averages input token embeddings—requiring no forward pass—matches or outperforms these more computationally intensive models.
+Adding a forward pass in this specialized model yields further improvements.
+These results strongly suggest that embedding models for reasoning about similarity between pretraining examples should be specialized to the data distribution at hand, and that properties that make embeddings suitable for retrieval and semantic matching may not transfer to the pretraining setting.
+More broadly, our work establishes a standardized framework to evaluate and highlight failures of current off-the-shelf models, facilitating the design of new models tailored for data curation in language model pretraining.
 2 Related Work
@@
 One family of works approaches data curation with the goal of retaining only high quality examples.
-The simplest approaches in this family look at heuristics that can filter out noisy or extremely low-quality instances such as documents containing less than 5 sentences or those with extremely short lines of text (Raffel et al., 2020; Xue, 2020).
-Beyond these heuristic approaches, a line of work focus on extracting high-quality examples based on similarity to known high-quality sources, such as Wikipedia or textbooks (Gunasekar et al., 2023; Penedo et al., 2024).
+The simplest approaches in this family look at heuristics that can filter out noisy or extremely low-quality instances, such as documents containing less than 5 sentences or those with extremely short lines of text (Raffel et al., 2020; Xue, 2020).
+Beyond these heuristic approaches, a line of work focuses on extracting high-quality examples based on similarity to known high-quality sources, such as Wikipedia or textbooks (Gunasekar et al., 2023; Penedo et al., 2024).
 Other works look at creating targeted sets—defining quality in terms of their relevance for particular downstream tasks (Xie et al., 2023), with some creating these by adaptively updating weights over domain mixtures (Xie et al., 2024; Jiang et al., 2024).
-The other family of data curation approaches focus on pruning datasets to ensure a notion of diversity (Abbas et al., 2023, 2024; Tirumala et al., 2023).
-While some of these works (Abbas et al., 2024; Tirumala et al., 2023) compare different embedding models as part of brief ablations, these comparisons are in the narrow context of their specific divsersification algorithms, and do not provide intuition as to why one embedding may perform better than another.
-Finally, it is worth mentioning that some works in data selection use the model being trained as part of the selection process, often through the use of influence functions or gradient-based methods (Garima et al., 2020; Xia et al., 2024; Engstrom et al., 2024).
-However, these are typically used during for data-curation for multi-epoch training on small datasets or only during the finetuning stage, and would be prohibitively expensive to use for pretraining.
+The other family of data curation approaches focus on pruning datasets to ensure a notion of diversity and reducing redundancies in pretraining examples (Abbas et al., 2023, 2024; Tirumala et al., 2023).
+While some of these works (Abbas et al., 2024; Tirumala et al., 2023) compare different embedding models as part of brief ablations, these comparisons are in the narrow context of their specific diversification algorithms.
+Finally, it is worth mentioning that some works in data selection use the model being trained as part of the selection process, often through the use of influence functions (Garima et al., 2020; Xia et al., 2024; Engstrom et al., 2024).
+However, these are typically only used in data-curation on small datasets or only during finetuning, and would be prohibitively expensive in pretraining.
 Text Embedding Models.
@@
 Many various approaches to learn text embedding models, with objectives including mask-based reconstruction (Devlin, 2018; Liu et al., 2019), a combination of multiple different tasks (Cer, 2018), and contrastive learning-based approaches (Gao et al., 2021; Neelakantan et al., 2022; Izacard et al., 2022; Lee et al., 2024).
-More recent work has demonstrated extracting embeddings via prompting, specifically designed for predicting model performance (Sam et al., 2025).
-These embedding models have largely been studied in the context for classification or similarity measures (Gomaa et al., 2013; Agirre et al., 2013, 2016), and recent focus has been improving performance on aggregate on large-scale benchmarks (Muennighoff et al., 2022) that are comprised of multiple tasks (e.g., retrieval, clustering, classification, ranking).
+More recent work has studied extracting embeddings from standard decoder-only language models (BehnamGhader et al., 2024), even via prompting (Sam et al., 2025).
+These embedding models have largely been studied in the context for classification or similarity measures (Gomaa et al., 2013; Agirre et al., 2013, 2016), and recent focus has been on improving performance on aggregate on large-scale benchmarks (Muennighoff et al., 2022) that are comprised of multiple tasks (e.g., retrieval, clustering, classification, ranking).
 Many models that achieve strong performance on these large-scale benchmarks have benefitted from scaling (Jiang et al., 2023; Chowdhery et al., 2023), with the help of distilled information from autoregressive models (Lee et al., 2024).
@@
 To the best of our knowledge, our work provides the first study of various text embedding models for pretraining data curation.
-3 Evaluating Embeddings for Pretraining Data Curation
-We now describe our approach to analyze and evaluate a given embedding model in terms of its suitability for reasoning about similarity among pretraining examples.
-As mentioned above, similarity is used to (a) find examples that are similar to known “desirable” examples (e.g., examples of known high-quality or those representing downstream tasks of interest) and to (b) discover and remove redundancies in the pretraining corpus.
-Accordingly, we design experiments that measure an text embedding model’s performance towards these criteria.
+3 An Evaluation Framework for Embeddings in Pretraining Data Curation
+We now describe our new framework for analyzing text embedding models in terms of their suitability for reasoning about similarity among pretraining examples.
+As previously mentioned, similarity is used to (a) find examples that are similar to known “desirable” examples (e.g., examples of known high-quality or those representing downstream tasks of interest) and to (b) discover and remove redundancies in the pretraining corpus.
+Accordingly, we design standardized experiments that measure a text embedding model’s performance towards these criteria.
 3.1 Evaluating Correlation with Pretraining Loss
-We begin by evaluating whether low distances in a model’s embedding space correlates with similar values of difficulty, which we measure by intermediate losses during language model pretraining.
-To do so, we first use a balanced K-Means clustering algorithm in the given embedding space to cluster all examples in the pretraining set for a target cluster size.
+We begin by evaluating whether low distances in a model’s embedding space correlate with similar values of difficulty, which we measure by intermediate losses during language model pretraining.
+Examples of positive correlations are visualized in Figure˜1.
+To do so, we first use a balanced K-Means clustering algorithm in the embedding space to cluster all examples in the pretraining set for a target cluster size.
 Then, we look at the variance of loss values within each cluster.
-We note that we use balanced clustering (allowing only some variation in the sizes of different clusters) to ensure that average within-cluster variance is comparable across clusterings in different embedding spaces.
-We repeat this process for a multiple target cluster sizes for all embeddings.
+This within-cluster variance captures whether points with similar pretraining loss (or difficulty) are grouped together and are close in embedding space.
+Crucially, we note that the converse is not necessarily true; two very different examples can have the same loss because they happen to be equally difficult to learn.
+This asymmetry is an important reason we chose our cluster-based approach rather than alternatives such as pairwise correlation.
+We note that we use balanced clustering (allowing only some variation in the sizes of different clusters) to ensure that average within-cluster variance is comparable across clusterings from different embedding models.
+We repeat this process for multiple target cluster sizes for all embeddings.
 When reporting our results, we measure variance reduction, or the ratio of the overall variance across all examples in the pretraining dataset to the variance in loss computed within clusters in the given embedding space.
-Specifically, we define variance reduction as
+Formally, we define variance reduction as
 $V(C)$ $=\frac{E_{x\sim D}[(\ell(x)-E[\ell(x)])^{2}]}{E_{C_{i}\sim\mathbbm{C}}\left[E_{x\sim C_{i}}[(\ell(x)-E_{x\sim C_{i}}[\ell(X)])^{2}]\right]},$
-where $C=\{C_{1},...,C_{m}\}$ is a partition of all datapoints in the dataset $D$, and $\mathbbm{C}$ is a uniform distribution over each disjoint set in the partition $C$.
-This metric can be interpreted as a measure of predicting pretraining generalization, as nearby points have similar loss values.
-We remark that random clusters achieve a variance reduction of 1, and larger values of variance reduction imply that points with more similar pretraining loss are clustered together.
+where $C=\{C_{1},...,C_{m}\}$ is a clustering or partition of all datapoints in the dataset $D$, and $\mathbbm{C}$ is a uniform distribution over each disjoint set in the partition $C$.
+Random clustering achieves a variance reduction of 1, and larger values of variance reduction imply that points with more similar pretraining loss are clustered together.
 A high value of variance reduction implies that similarity in the embedding space correlates well with pretraining loss.
-While an isolated pair of examples having similar pretraining loss may be entirely unrelated in quality and content, the fact that an embedding space that consistently brings together examples of similar loss values strongly suggests that these similar examples will behave similarly in terms of contributing to the quality of the language model.
+While an isolated pair of examples having similar pretraining loss may be entirely unrelated in quality and content, the fact that an embedding space that consistently brings together examples of similar loss values strongly suggests that these similar examples will behave similarly in terms of contributing to the quality of the language model (when used in pretraining).
 Beyond finding examples of similar quality and utility, a high variance reduction score implies that the clustering above can serve as a particularly useful proxy for more dynamic and online data sampling strategies (Xie et al., 2024), reducing the number of required forward passes in strategies where datapoints are selected based on their current pretraining loss.
 This also has implications towards approaches that look to self-improve models through propagating labeled information to nearby examples (Wei et al., 2021; Cai et al., 2021; Pukdee et al., 2023), such as weak-to-strong generalization (Burns et al., 2023).
-3.2 Measuring Cluster Purity of Data Sources
-Next, we study the purity of the respective (K-Means) clusters generated with each considered embedding model with respect to the underlying data source or mixture.
+3.2 Diversification-based Pretraining Data Curation
+Prior work has demonstrated that data curation schemes that encourage diversification, based on similarity in embedding spaces, leads to improved models when trained on the curated subsets (Abbas et al., 2023; Tirumala et al., 2023).
+While the diversification techniques tend to be sophisticated, our goal here is to evaluate the utility of a given embedding for diversification.
+Therefore, we use a simplified version of these approaches to (1) select a subset from a pretraining corpus, (2) pretrain a model on this curated subset, and then (3) report performance on a large set of downstream tasks.
+We begin by clustering the larger pretraining corpus in the given embedding space; in contrast to Section˜3.1, we do not use balanced clustering, which can lead to clusters with a wide variation in “diameters” (distances among points in the same cluster).
+Instead, since the goal here is diversification, we find clusters such that distances between all pairs of points in the same cluster are within a specified threshold $\epsilon$.
+Given clusters of pretraining data, prior work often uses complex pruning or sampling schemes from large clusters.
+For simplicity, we simply take the point in each cluster that is closest to the cluster centroid (i.e., the average embedding of all points in the cluster), which likely represents the most representative example for that cluster.
+We note that prior work (Tirumala et al., 2023) emphasizes the importance of a deduplication step in language model pretraining, and our curation strategy implicitly performs this process when selecting only a single point from each cluster.
+Given that we select a single point from each cluster, this implies that we need the number of clusters to be equal to or greater than the number of desired pretraining examples.
+In practice, we sweep a large grid of $\epsilon$ values and select the largest threshold that still produces a sufficient number of clusters.
+Another important point of note is that producing such a large number of clusters is computationally expensive.
+To be able to scale to such a large number of clusters, we use the reciprocal agglomerative clustering (RAC) (Sumengen et al., 2021) algorithm.
+This has the advantage of building a clustering from progressively larger values of $\epsilon$, which works well with our need for a grid of values of $\epsilon$.
+Finally, we pretrain a language model on the selected subset of examples, and measure few-shot performance of the pretrained model on a diverse variety of downstream tasks.
+3.3 Measuring Cluster Purity with Respect to Data Sources
 Large pretraining datasets such as the Pile (Gao et al., 2020) are comprised of various distinct yet complementary hand-curated data sources (e.g., high quality sources such as Wikipedia, code-based data, data from medical domains).
-While there may be some similarities between datapoints from different domains, we believe that clusters should generally group points from the same domain together and separate isntances from different domains.
-In fact, we remark that since embedding models are generally not trained with explicit knowledge of source labels or any domain metadata, and the ability to group together data from the same source and distinguish those from different sources shows an embedding model’s alignment with human judgement (or at least that of the dataset curators).
+While there may be some similarities between datapoints from different domains, we believe that clusters should generally group points from the same domain together and separate instances from different domains.
+In fact, since embedding models are generally not trained with explicit knowledge of source labels or any domain metadata, the ability to group together data from the same source and distinguish those from different sources shows an embedding model’s alignment with meaningful human-curated structure (or at least that of the dataset curators).
+This also serves as a useful proxy of embedding model capacity, as they should contain sufficient information to well-separate these different sources.
 Letting $s\in\{0,\cdots,S\}$ represent the index of the source in some finite number of sources $S$, we compute the cluster purity of a set of clusters $C$ as
@@
 A cluster purity of 1 represents a clustering that completely separates examples from different domains.
-3.3 Diversification-based Pretraining Data Curation
-Prior work has demonstrated that diversification, based on similarity in embedding spaces, can yield more effective pretraining datasets that in-turn lead to improved models (Tirumala et al., 2023).
-While the diversification techniques themselves tend to be sophisticated, our goal here is to evaluate the utility of a given embedding for diversification.
-Therefore, we use a simplified version of these diversification of approaches to (1) select a subset from a larger pretraining corpus, (2) pretrain a model on this curated subset, and then (3) report performance on a large set of downstream tasks.
-Figure 2: Reduction in variance in loss among clusters (e.g., overall variance / clustering variance) as increase the number of gradient steps in pretraining (left).
-Reduction in variance as we vary the approximate cluster size used in K-means for clustering examples with various embedding models (right).
-We compute our results on the checkpoint after 26k gradient steps.
-In both figures, we compute our results over 50 million sampled clusters from the Pile, and we remark that larger values are better.
-The overall pretraining loss variances for checkpoints of 26k, 50k, and 76k training steps are 0.521, 0.508, and 0.489 respectively.
-We find that embeddings extracted from a small version of the downstream language model being trained best cluster points with similar pretraining losses, across different intermediate training checkpoints and cluster sizes.
-We begin by clustering the larger pretraining corpus in the given embedding space; in contrast to Section 3.1 and Section 3.2, we do not use balanced clustering which can lead to clusters with a wide variation in “diameters” (distances among points in the same cluster).
-Instead, since the goal here is diversification, we find clusters such that distances between all pairs of points in the same cluster are within a specified threshold $\epsilon$.
-Given clusters of pretraining data, prior work often uses complex pruning or sampling schemes from large clusters.
-For simplicity, we simply take the point in each cluster that is closest to the cluster centroid (i.e., the average embedding of all points in the cluster), which likely represents the most representative example for that cluster.
-We note that prior work (Tirumala et al., 2023) emphasizes the importance of a deduplication step in language model pretraining, and our curation strategy implicitly performs this process when selecting only a single point from each cluster.
-Given that we select a single point from each cluster, this implies that we need the number of clusters to be equal or greater to the number of desired pretraining examples.
-In practice, we sweep a large grid of $\epsilon$ values and select the largest threshold that still produces a sufficient number of clusters.
-Another important point of note is that producing such a large number of clusters is computationally expensive.
-To be able to scale to such a large number of clusters, we use the reciprocal agglomerative clustering (RAC) (Sumengen et al., 2021) algorithm.
-This has the advantage of building a clustering from progressively larger values of $\epsilon$, which works well with our need of a grid search over values of $\epsilon$.
-Finally, we pretrain a language model on the selected subset of examples, and measure few-shot performance of the pretrained model on a diverse variety of downstream tasks.
 4 Experiments
-We analyze various different approaches to extract text embeddings across multiple different pretraining settings.
+We demonstrate the utility of our framework in evaluating off-the-shelf embedding models in measuring pretraining similarity and compare them with simple ways to produce specialized embeddings.
 4.1 Experiment Details
-Text Embedding Models
-For text embedding models, we compare Universal Sentence Encoders (USE) (Cer, 2018), which are general purpose text encoders, and Gecko (Lee et al., 2024), a retrieval-focused text embedding model trained via synthetic data distilled from LLMs, and BERT (Devlin, 2018) models.
-We also compare against various techniques to extract embeddings from a small version of the downstream language model we are training (e.g., a  200M parameter language model).
-First, we extract an embedding by simply averaging the token embeddings matrix over all tokens in the input sequence, which we refer to as LM Token Embeds.
-We note that this is equivalent to a learned unigram model (i.e., ignoring all positional information).
-We also extract an embeddings from a forward pass of the language model by averaging the activations over all tokens in the input sequence at the final hidden layer, which we refer to as LM Output Embeds.
-In our diversity-based pretraining data curation experiments, we also add a comparison to a naive, random subset selection (Baseline) to evaluate how much benefit is added from our standardized curation strategy with these various embedding models.
+Embedding Models
+For the off-the-shelf embedding models, we consider: (1) Universal Sentence Encoders (USE) (Cer, 2018), which are general-purpose text encoders trained on a combination of objectives, (2) Gecko (Lee et al., 2024), a retrieval-focused text embedding model trained via synthetic data distilled from LLMs, and (3) BERT (Devlin, 2018) embedding models.
+These encapsulate a variety of different training objectives and reflect common embedding model choices in the field.
+We also consider a few standard approaches to extract specialized embeddings that are extracted from a small version of the downstream language model we are training (e.g., a  200M parameter language model).
+The first and most simple approach is to extract an embedding by simply averaging the token embeddings matrix over all tokens in the input sequence, which we refer to as LM Token Embeds.
+This is extremely efficient as it only requires looking up the token embedding matrix and does not involve any forward passes.
+We note that this is also equivalent to a learned unigram model (i.e., ignoring all positional information).
+The second approach is to extract an embedding from the forward pass of the language model by averaging the activations over all tokens in the input sequence at the final hidden layer, which we refer to as LM Output Embeds.
+Finally, in our diversity-based pretraining data curation experiments, we also add a comparison to a naive, random subset selection (Baseline).
+This evaluates how much benefit is observed from our standardized curation strategy with these various embedding models over standard baseline training.
+Clustering Details
 To make clustering with a large number of output clusters feasible and efficient at pretraining data scales, we perform dimensionality reduction using PCA on a large subset of the datapoints ($\sim$500,000 examples) and extract the top 64 dimensions upon which to project.
 The projected embeddings are then normalized to have a unit L2 norm.
-We also explore using random projections (Bingham & Mannila, 2001) as a different dimensionality reduction strategy, finding that random projections do not perform as well as PCA (see Section 4.5 for ablation studies).
-Dataset Details
+We ablate on this choice for dimensionality reductions by comparing with random projections (Bingham and Mannila, 2001) as an alternative dimensionality reduction strategy; we find that random projections do not perform as well as PCA (Section˜4.5).
+Other work (Tirumala et al., 2023) tackles this computational challenge by running clustering only with a much smaller subset of data, rather than performing dimensionality reduction and then clustering over all the data.
+Figure 2: Variance reduction as we vary average cluster size.
+Larger values are better.
+Results are computed over 50 million sampled clusters from the Pile, where pretraining losses are computed after 26k gradient steps.
+Specialized embeddings yield higher variance reduction than off-the-shelf models for all cluster sizes.
+Figure 3: Variance reduction as we increase the number of gradient steps in pretraining.
+Larger values are better.
+Results are computed over 50 million sampled clusters from the Pile with an average cluster size of 50.
+Benefits in variance reduction remain consistent throughout pretraining.
+Dataset and Pretraining Details
 For all of our experiments (e.g., pretraining data curation and predicting loss generalization), we use the Pile (Gao et al., 2020).
-We pack together documents into a sequence of length 1280, with “[eod]” as delimeters between documents.
-Pretraining Details
-For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022), on the curated subset of the Pile data through the process outlined in Section 3.3.
-We use a selection budget of 200B tokens, which is approximately 20% of sequences from the Pile.
+We pack together documents into a sequence of length 1280, with “[eod]” as delimiters between documents.
+For our pretraining experiments, we train a 1.7B parameter decoder-only language model with a UL2 objective (Tay et al., 2022).
+We curated pretraining subsets from the Pile through the process outlined in Section˜3.2.
+We use a selection budget of 200B tokens, or approximately 20% of the Pile.
 This corresponds to roughly 170 million clusters for each embedding model, with an average cluster size of 5 examples.
@@
 For our tokenizer, we use sentencepiece with a vocabulary size of 256k tokens.
-For our pretraining evaluation, we consider the same downstream evaluation datasets from the work of Brown et al.
+For our pretraining evaluation, we consider the set of 23 downstream evaluation datasets from the work of Brown et al.
 (2020).
-This invovles a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks.
-The full details about our evaluation sets are deferred to Section B.1.
-For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we train on a total of 200B randomly selected tokens.
+This involves a wide variety of 1-shot scoring tasks, as well as open-ended text generation tasks.
+More details about our evaluation sets are deferred to Section˜B.1.
+For the embeddings produced via LM Token Embeds and LM Output Embeds, we train a 200M parameter language model in the same fashion as above, where we also train on a total of 200B randomly selected tokens.
+Table 1: Average downstream task performance of embedding models in the diversity-based curation schemes of a 200B token subset from the Pile, to pretrain a 1.7B parameter decoder-only language model.
+Bolding and italicizing denote the best and second-best performing methods on each task, respectively.
+Results are averaged over 3 pretraining runs, and average results are mean $\pm$ standard error.
+Specialized embeddings (i.e., extracted from a small version of the downstream language model) and Gecko perform the best for diversity-based pretraining data curation.
+Task Baseline USE Gecko BERT LM Token Embeds LM Output Embeds
+ARC - Challenge 32.4 33.0 33.7 32.7 32.5 33.5
+ARC - Easy 63.8 65.1 64.2 65.1 64.9 65.5
+BoolQ 56.5 59.2 58.3 60.1 62.9 61.9
+SuperGLUE - CB 42.4 43.5 48.2 41.7 42.9 48.8
+SuperGLUE - Copa 75.3 76.3 77.3 74.0 74.7 78.0
+HellaSwag 55.0 56.7 57.1 56.5 56.8 57.5
+Multi RC 57.7 55.7 55.9 56.3 56.0 55.2
+OpenBook QA 46.3 46.2 46.3 46.6 45.9 46.7
+PiQA 72.3 73.0 72.7 73.6 72.8 73.6
+Race H 38.0 37.9 38.5 38.8 38.8 38.7
+Race M 51.7 52.0 53.3 52.9 52.1 52.6
+ReCoRD 85.0 84.4 84.6 84.7 84.5 84.9
+RTE 54.5 54.9 51.5 54.9 51.6 52.8
+Story Cloze 73.9 74.1 74.3 74.0 73.3 74.5
+WiC 48.2 47.9 47.5 47.5 48.2 47.3
+Winograd 74.0 76.0 77.1 77.3 77.1 76.6
+WinoGrande 59.2 58.9 59.0 59.5 60.0 59.3
+WSC 74.1 73.9 73.7 74.7 73.2 75.0
+Lambada 21.4 29.9 40.1 34.7 31.4 33.8
+Natural Questions 10.0 10.8 10.1 10.5 11.1 11.4
+SQuAD v2 51.1 54.3 51.7 51.8 58.6 55.4
+TriviaQA Wiki 34.4 35.0 33.0 35.0 34.3 36.1
+Web Questions 17.1 17.0 17.9 18.0 18.7 19.1
+Average 51.9 $\pm$ 0.1 52.9 $\pm$0.1 53.3 $\pm$ 0.3 53.1 $\pm$ 0.2 53.1 $\pm$ 0.3 53.8 $\pm$ 0.1
 4.2 Pretraining Loss Correlation Results
-To measure the ability of embedding models to reflect pretraining loss generalization and to cluster together datapoints with similar difficulty, we report their variance reduction as we vary (1) the number of gradient updates in language model pretraining and (2) as we vary the average cluster size in our K-Means clustering.
-We observe that embeddings produced via a small UL2 model (i.e., LM Token Embeds and LM Output Embeds) trained on the same data achieve a much larger reduction in loss variance when compared to off-the-shelf models; these embeddings produce better clusters with datapoints with similar intermediate pretraining losses (Figure 2).
-These trends hold across multiple different intermediate training checkpoints and for all different cluster sizes.
+To measure the ability of embedding models to reflect pretraining loss generalization and to cluster together datapoints with similar difficulty, we report their variance reduction as we vary (1) the number of gradient updates performed by the intermediate pretraining checkpoint and (2) as we vary the average cluster size in our K-Means clustering.
+We observe that specialized embeddings (i.e., LM Token Embeds and LM Output Embeds) from a small version of the downstream trained model, which is trained on the same pretraining data, achieve a much larger reduction in loss variance when compared to off-the-shelf models.
+These specialized embeddings produce better clusters with datapoints with similar intermediate pretraining losses, across all cluster sizes (Figure˜3) and for all intermediate pretraining checkpoints (Figure˜3).
 This suggests that training models specifically on the data to perform data curation is preferable to using off-the-shelf embedding models for predicting pretraining loss generalization.
-Furthermore, we also generally observe that newer embedding models (e.g., Gecko) do not improve upon older alternatives (e.g., BERT and USE), supporting that embedding model improvements in other settings such as retrieval do not translate to this application in the pretraining setting.
+Furthermore, we also generally observe that newer embedding models (e.g., Gecko) do not improve upon older alternatives (e.g., BERT and USE), supporting that embedding model improvements in other settings, such as retrieval do not translate to this application in the pretraining setting.
 Visualizing Within-Cluster Examples
-To better interpret what is computed by these similarity metrics, we also visualize pairs of datapoints from the various clusters in Figure 1 to understand what types of examples are placed closely together and which have similar pretraining losses.
-We randomly sample 3 clusters produced by K-Means clusters with the smallest average cluster size with the LM Output Embeds method to extract embeddings.
-We randomly sample 2 points from these 3 clusters and observe that samples from the same cluster are thematically similar, although they contains very different information (e.g., the last example referring to two different people in history with some governmental relations, although from different times and different countries).
-We observe that pretraining losses are similar between within-cluster examples, and lower distances correspond to lower differences in pretraining loss.
+To better interpret these similarity metrics, we visualize pairs of datapoints from the clusters in Figure˜1 to understand what types of examples are placed closely together and which have similar pretraining losses.
+We randomly sample 3 clusters produced by K-Means clusters with the smallest average cluster size, with the LM Output Embeds method to extract embeddings.
+We randomly sample 2 points from these 3 clusters and observe that samples from the same cluster are thematically similar, although containing very different information (e.g., the last pair referring to two different government-related people, although from different times and different countries).
+We observe that pretraining losses are similar among within-cluster examples, and lower distances correspond to smaller differences in pretraining loss.
+4.3 Diversification-based Pretraining Data Curation Results
+We report the average downstream task performance of language models trained on pretraining data subsets produced by using various embedding models in the simple diversity-based data curation strategy in Table˜1.
+We first observe that the standardized diversity-based data curation technique applied to all embedding models (both off-the-shelf and those based on small downstream trained language models) outperforms the naive random subset selection baseline.
+Secondly, we observe that the LM Output Embeds achieves the best performance when averaged across all tasks.
+A notable result is that LM Token Embeds performs comparably to Gecko, while outperforming all other off-the-shelf embeddings.
+This is of immediate practical interest due to the simplicity of LM Token Embeds; they require no forward pass and, consequently, are extremely quick to compute in comparison to all other considered alternatives.
+Thus, LM Token Embeds serves as a viable alternative in settings that are compute-limited.
+This also suggests that even simple notions of similarity, even those that do not account for positional information, are sufficient for many pretraining data curation applications.
+Finally, they support that models specialized to this task (e.g., trained on the same dataset) often outperform the current, general-purpose, off-the-shelf embedding models.
+While this does include training an additional model, it is a single-time investment that is amortized across all embedding use cases and is easily offset by the significant improvements in data curation quality.
 4.4 Cluster Purity Results
-Figure 3: Comparison of the purity with respect to data source of K-Means clustering produced by various embedding models on the Pile. Cluster purities are averaged over 50 million clusters.
+Figure 4: Comparison of the purity with respect to data source of K-Means clustering produced by various embedding models on the Pile, when averaged over 50 million clusters from the Pile.
+Specialized embedding models have higher cluster purity scores.
 To better understand and interpret the similarity metrics defined by these embedding models, we measure the purity of clusters with respect to the underlying data source.
 We again remark that none of these embedding models have been trained with knowledge of the data source.
-Overall, we observe that most embedding models produce fairly pure clusters, where a majority of points come from the same underlying data source (Figure 3).
-This supports that these embeddings models are generally aligned with human judgement about differences between types of data.
+Overall, we observe that most embedding models produce fairly pure clusters, where a majority of points come from the same underlying data source (Figure˜4).
+This supports that these embeddings models are generally aligned with human judgment about differences between types of data.
 We observe that the embeddings extracted from the small version of the trained downstream language model achieve the highest cluster purity.
@@
 This suggests that an improved ability to predict pretraining loss generalization cannot simply be explained via producing clusters that are more pure with respect to underlying data sources.
-Future work could explore incorporating domain information or metadata into the embedding model’s learning objective, as a form of weak supervision (Sam & Kolter, 2023).
-Task Baseline USE Gecko BERT LM Token Embeds LM Output Embeds
-ARC - Challenge 32.44 32.99 33.73 32.71 32.54 33.53
-ARC - Easy 63.78 65.12 64.24 65.08 64.91 65.46
-BoolQ 56.47 59.18 58.31 60.08 62.86 61.93
-SuperGLUE - CB 42.41 43.45 48.21 41.67 42.86 48.81
-SuperGLUE - Copa 75.25 76.33 77.33 74.00 74.67 78.00
-HellaSwag 54.98 56.71 57.11 56.54 56.76 57.53
-Multi RC 57.71 55.69 55.91 56.27 55.95 55.19
-OpenBook QA 46.30 46.20 46.27 46.60 45.93 46.67
-PiQA 72.27 72.98 72.74 73.58 72.76 73.63
-Race H 38.02 37.94 38.47 38.75 38.84 38.70
-Race M 51.71 52.00 53.32 52.88 52.09 52.60
-ReCoRD 84.95 84.35 84.64 84.71 84.51 84.91
-RTE 54.51 54.87 51.50 54.87 51.62 52.83
-Story Cloze 73.93 74.10 74.26 73.95 73.26 74.49
-WiC 48.20 47.86 47.49 47.54 48.17 47.28
-Winograd 73.99 75.95 77.05 77.29 77.05 76.56
-WinoGrande 59.23 58.91 59.01 59.46 59.96 59.25
-WSC 74.12 73.92 73.68 74.74 73.22 74.97
-Lambada 21.37 29.86 40.12 34.69 31.43 33.82
-Natural Questions 10.01 10.76 10.13 10.54 11.13 11.42
-SQuAD v2 51.05 54.26 51.70 51.79 58.56 55.37
-TriviaQA Wiki 34.39 35.01 32.99 34.95 34.29 36.08
-Web Questions 17.05 17.04 17.87 17.98 18.73 19.13
-Average 51.92 52.85 53.30 53.07 53.13 53.83
-Table 1: Performance of using various embedding models in the diversity-based selection scheme to select a subset of 200B tokens from the Pile to pretrain a 1.7B parameter language model with a UL2 (Tay et al., 2022) objective.
-Bolding and underlining denote the best and second-best performing methods on each task, respectively.
-Results are averaged over 3 pretraining runs.
-We find that embeddings extracted from small version of the downstream language model and Gecko perform the best for pretraining data curation.
-4.4 Diversification-based Pretraining Data Curation Results
-We report the average downstream task performance of language models trained on pretraining data subsets produced by using various embedding models in the simple diversity-based data curation strategy in Table 1.
-We first observe that the standardized diversity-based data curation technique applied to all embedding models (both off-the-shelf and those based on small downstream trained language models) outperform the naive random subset selection baseline.
-Secondly, we observe the LM Output Embeds and Gecko embeddings achieve the best and second-best performance when averaged across all tasks, respectively.
-A notable result is that LM Token Embeds performs comparably to Gecko, while outperforming all the other off-the-shelf embedding models.
-This is of practical interest due to the simplicity of producing these embeddings; they are simply constructed via averaging over the learned token embeddings and, consequently, are extremely quick to compute in comparison to all other embedding methods.
-Thus, LM Token Embeds serves as a powerful alternative in settings that are compute-limited.
-These pretraining results also suggest that even simple notions of similarity, even those that do not account for positional information, are sufficient for many pretraining data curation applications.
-Finally, they support that the training of models specifically for this task (e.g., on the same pretraining dataset) is preferable to the use of general purpose off-the-shelf embedding models.
+Future work could explore incorporating domain information or metadata into the embedding model’s learning objective, as a form of weak supervision (Sam and Kolter, 2023).
 Figure 5: Ablation on the number of components in PCA for Gecko and LM Output Embeds.
 Results are averaged over 50 million sampled clusters from the Pile.
-Using more components in PCA better cluster points with similar pretraining loss.
+Using more components in PCA better clusters points with similar pretraining loss.
 Figure 6: Ablation comparing the use of PCA or Random projections for dimensionality reduction in Gecko and LM Output Embeds.
 Results are averaged over 50 million sampled clusters from the Pile.
-Dimensionality reduction via PCA performs better for variance reduction when compared to Random Projections.
+Dimensionality reduction via PCA performs better than via Random Projections.
 4.5 Ablations
-We present ablations on the various dimensionality reduction strategies (i.e., technique and resulting size) that we use for the embedding models before running clustering.
-We present the ablations on Gecko and LM Output Embeds here and defer results on others embeddings to Appendix A.
-Impact of Dimensionality on Variance Reduction
+We present ablations on components of our evaluation framework — specifically in how we have chosen to perform dimensionality reduction (i.e., technique and resulting size), which is used for the embedding models before running clustering.
+We present the ablations for Gecko and LM Output Embeds and defer results on others embeddings to Appendix A.
+Less Components in PCA Does Not Significantly Hurt Performance.
 The dimensionality of the embeddings used in clustering often must be low for pretraining scales.
-Here, we run an ablation studying the role of dimensionality (i.e., the number of components in PCA onto which the embeddings are projected) in the variance reduction through clustering (Figure 4).
+Here, we run an ablation studying the role of dimensionality (i.e., the number of components in PCA onto which the embeddings are projected) in the variance reduction through clustering (Figure˜5).
 We remark that scaling clustering to accommodate smaller average cluster sizes (i.e., more cluster centers) is intractable for embeddings that have high dimensionality.
@@
 We observe the trend across all embeddings that a higher dimension and larger number of PCA components improves the embedding models’ ability to cluster points by pretraining loss (see other embedding model results in Appendix A.1).
-Dimensionality Reduction Techniques
+Dimensionality Reduction with PCA Outperforms Random Projections.
 Another common technique to perform dimensionality reduction (especially with respect to maintaining pairwise distances and similarities) is to use random projections.
-We experiment with using random projections in our loss clustering experiments and observe that it is outperformed by PCA in terms of variance reduction across all cluster sizes (Figure 5).
+We experiment with using random projections in our loss clustering experiments and observe that it is outperformed by PCA in terms of variance reduction across all cluster sizes (Figure˜6).
 5 Discussion
-We analyze the fundamental role of computing similarity in pretraining data curation, finding that coarse-grained notions of similarity are particularly useful and that specific models (even simple ones based on averaging learned token embeddings) can outperform other off-the-shelf embedding models.
-We believe that both our results and the new evaluation framework for similarity metrics provide new insights into the role of similarity in pretraining data curation.
-For instance, our variance reduction results shows that embedding models tightly cluster points with similar values of pretraining losses, which has larger implications for dynamic data selection.
-Future work could study that, during training, we can compute the loss on the centroid in each cluster, and use this to select clusters on which to train, that contain points with high values of these intermediate loss.
-Our method has other, broad potential applications in the pretraining data setting, such as selecting task-specific data for finetuning (Xia et al., 2024) or even in evaluating the generation of synthetic data for (Meng et al., 2022; Sam et al., 2024) by looking at the distance between these potential data to be selected or generated and real data.
-Finally, our results highlight the benefits for designing models specifically for language model pretraining settings, and our framework provides an evaluation to exactly capture their utility.
+We present a new evaluation framework for the understudied area of similarity metrics used in language model pretraining data curation.
+Using our framework, we show that off-the-shelf embeddings—despite their widespread use—often underperform even simple, specialized alternatives such as an average of token embeddings that requires no forward pass.
+This also suggests that practitioners should train their own embedding models, and it suffices to train them on a small fraction of the data ($\sim$20%) and at a much smaller parameter scale.
+While it may seem intuitive that embeddings tailored to the pretraining task would outperform generic ones, this has not been systematically demonstrated and is certainly not yet standard practice.
+Our framework both surfaces this gap and offers a practical tool for guiding the design of embedding models optimized for data curation, which can lead to significant improvements in data efficiency.
+While we believe that our results are fairly general, future work could extend our findings to other pretraining corpora.
+Beyond evaluation, our findings have broader implications for pretraining workflows—for example, in selecting task-specific finetuning data (Xia et al., 2024), or identifying synthetic examples that resemble real data (Meng et al., 2022; Sam et al., 2024), and even studying the impacts of scaling up embedding models in these tasks.
+Overall, our results underscore the importance of using task-aligned similarity metrics in pretraining, and our framework provides a foundation for future research in building and optimizing the design of embedding models tailored to this critical step of the language modeling pipeline.
+Limitations
+One limitation of our work is that the empirical demonstration of our new evaluation framework is restricted to the Pile dataset.
+We believe this is fairly standard in the language model pretraining data curation literature to focus on a single pretraining corpus (e.g., as is done in Tirumala et al.
+(2023)), due to the computational costs of both data curation algorithms and pretraining.
 Appendix A Additional Experimental Results
@@
 B.1 Additional Evaluation Details
-We report results averaged over a large numer of downstream tasks.
+We report results averaged over a large number of downstream tasks.
 These datasets largely follow two categories: scoring and decoding tasks.
-Scoring tasks primarly look at the output distribution of the model, while decoding tasks look at text generations from the language model.
+Scoring tasks primarily look at the output distribution of the model, while decoding tasks look at text generations from the language model.
 Scoring tasks are performed as 1-shot (i.e., giving one demonstration of format and answer), while decoding is performed zeroshot.
@@
 The list of scoring tasks is as follows:
-  • ARC Challenge: Easy and Challenge (Clark et al., 2018)
-  • BoolQ (Clark et al., 2019)
-  • SuperGLUE - CB and Copa (Wang et al., 2019)
-  • HellaSwag (Zellers et al., 2019)
-  • MultiRC (Khashabi et al., 2018)
-  • OpenBookQA (Mihaylov et al., 2018)
-  • PIQA (Bisk et al., 2020)
-  • RACE-H and RACE-M (Lai et al., 2017)
-  • ReCoRD (Zhang et al., 2018)
-  • RTE (Dagan et al., 2010)
-  • Story Cloze (Mostafazadeh et al., 2016)
-  • WIC (Pilehvar & Camacho-Collados, 2018)
-  • Winograd (Levesque et al., 2012)
-  • WinoGrande (Sakaguchi et al., 2021)
-  • WSC (Levesque et al., 2012)
+  • ARC Challenge: Easy and Challenge [Clark et al., 2018]
+  • BoolQ [Clark et al., 2019]
+  • SuperGLUE - CB and Copa [Wang et al., 2019]
+  • HellaSwag [Zellers et al., 2019]
+  • MultiRC [Khashabi et al., 2018]
+  • OpenBookQA [Mihaylov et al., 2018]
+  • PIQA [Bisk et al., 2020]
+  • RACE-H and RACE-M [Lai et al., 2017]
+  • ReCoRD [Zhang et al., 2018]
+  • RTE [Dagan et al., 2010]
+  • Story Cloze [Mostafazadeh et al., 2016]
+  • WIC [Pilehvar and Camacho-Collados, 2018]
+  • Winograd [Levesque et al., 2012]
+  • WinoGrande [Sakaguchi et al., 2021]
+  • WSC [Levesque et al., 2012]
 For decoding or text generation tasks, we evaluate the language model outputs with its F1 score. Decoder tasks are also evaluated as 1-shot. The list of decoding tasks is as follows:
-  • Lambada (Paperno et al., 2016)
-  • Natural Questions (Kwiatkowski et al., 2019)
-  • Squad v2 (Rajpurkar et al., 2018)
-  • Trivia QA Wiki (Joshi et al., 2017)
-  • Web Questions (Berant et al., 2013)
+  • Lambada [Paperno et al., 2016]
+  • Natural Questions [Kwiatkowski et al., 2019]
+  • Squad v2 [Rajpurkar et al., 2018]
+  • Trivia QA Wiki [Joshi et al., 2017]
+  • Web Questions [Berant et al., 2013]
 B.2 Additional Embedding Model Details
@@
 For our K-Means clustering, we perform clustering at 4 different levels of granularity in our variance reduction and cluster purity results.
-We create four sets of clusterings with an average cluster size of 25, 50, 100, 150, with a minimum cluster size of $\frac{1}{5}$ times the average cluster size and a maximum cluster size of $5$ times the average cluster size.
+We create four sets of clusterings with an average cluster size of 25, 50, 100, 150, with a minimum cluster size of $\frac{1}{5}$ times the average cluster size, and a maximum cluster size of $5$ times the average cluster size.
 For both RAC and K-Means, we use the squared L2 (Euclidean) distance.
@@
 As previously mentioned, after projecting onto the desired number of principal components, we perform L2 normalization.
-For our random projections, we use a sparse random projections onto values of $-\frac{\sqrt{n}}{\sqrt{64}},0,\frac{\sqrt{n}}{\sqrt{64}}$ with probabilities $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ respectively (i.e., the default parameters via scikit-learn).
+For our random projections, we use a sparse random projections onto values of $\frac{\sqrt{n}}{\sqrt{64}},0,\frac{\sqrt{n}}{\sqrt{64}}$ with probabilities $\frac{1}{4},\frac{1}{2},\frac{1}{4}$ respectively (i.e., the default parameters via scikit-learn).
 We also L2 normalize the result of random projections.
@@
 Pretraining experiments for our 1.7B parameter language models are run on 512 v5 TPUs, where each pretraining run takes approximately 3 days.
+Training our proxy 200M parameter model took less than 1 day on 64 v5 TPUs.
+Hierarchical clustering for pretraining requires approximately 1-2 days to run over the full pretraining corpus.
+B.6 Asset Licenses
+The existing assets that we use have the following licenses:
+  • RAC Clustering: MIT license
+  • The Pile: CC BY 4.0
+  • Evaluation Datasets: MIT License
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: All claims are supported by experimental and theoretical results.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: Limitations are discussed in the Discussion section.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: No theoretical results are included.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: All experimental details are provided.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [No]
+  24. Justification: No code is provided, but all experimental details and evaluation metrics are clearly defined.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: All experimental details are provided in the code in the Appendix.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Error bars are provided when applicable.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: Compute details are provided in the Appendix.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: The submission conforms with the code of ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: Impacts of the paper are discussed in the discussion section.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: No such risks are posed.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: All assets are properly credited, and licenses are mentioned.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: No new assets are released.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [Yes]
+  79. Justification: All usage of LLMs (in computing similarities and in pretraining) are clearly described in the paper.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.